OpenStudy (anonymous):

use the quadratic formula to solve the equation: -3x^2-4x+10=0.... can you please sho me the working

OpenStudy (anonymous):

\[ax ^{2}+bx+c=0\] \[(-b \pm \sqrt{b ^{2}-4ac})/2a\]

OpenStudy (anonymous):

just plug in the the numbers to the corresponding letters

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

-3=a -4=b 10=c plug those into the second formula i gave you to get the answer

OpenStudy (anonymous):

i did that already

OpenStudy (anonymous):

so what did you get

OpenStudy (anonymous):

i don't know what else to do

OpenStudy (anonymous):

ok, well, the answer is -2.6103, and 1.2769839

OpenStudy (anonymous):

how is that

OpenStudy (anonymous):

(4+sqrt136)/-6 and (4-sqrt136)/-6

OpenStudy (anonymous):

2x^2+9x=5 a=2 b=9 c=-5 x=-9+squareroot 9^2-4(2)(-5) all over 2(2) x=-9+squareroot 81+40 over 4 x=-9+squareroot 121 over 4 =9+11 over 4 1 solution=- 9+11 over 4=1/2 2 solution= -9-11 over 4=-20/4 =-5

OpenStudy (anonymous):

like that

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

work out the next 1 like that

OpenStudy (anonymous):

is anyone there

OpenStudy (anonymous):

if your asking about 2x^2+9x=5 then your answer looks good to me

OpenStudy (anonymous):

am talk about -3x^-4x+10=0

OpenStudy (anonymous):

 am stuck there